I'm going to type the word problem. Help me with this.
Mary's class took two tests last week. 80% of the class passed the math test, 90% of the class passed the English test, and 72% of the class passed both. What is the probability that a randomly selected student in Mary's class failed both tests? Express your answer as a percent rounded to the nearest tenth of a percent.

Answer by stanbon(75887)

Mary's class took two tests last week. 80% of the class passed the math test, 90% of the class passed the English test, and 72% of the class passed both. What is the probability that a randomly selected student in Mary's class failed both tests? Express your answer as a percent rounded to the nearest tenth of a percent.
--------
P(passed M or passed E) = P(M) + P(E) - P(M and E)
P(passed Math or Eng) = 0.8 + 0.9 - 0.72 = 0.98
-----
P(did not pass Math or Eng) = 1-0.98 = 0.02
------------------------------------------------
Cheers,
Stan H.